### v2

@@ 5 Improving Generalization
We report in Table 2 and Figure 8 the results of averaging in terms of generalization performance. We find that using LAWA or EMA on top of a learning rate schedule consistently improves over the baseline optimizer, achieving slightly better validation scores across all the considered workloads. We notably observe significant improvements on the WMT workload. As previously reported in Kaddour (2022) and Sanyal et al. (2023), in the early stage of training, averaging schemes provide considerably better performance than the underlying optimization algorithm, but this gap shrinks towards the end of training. We argue that this behavior is closely linked to the use of WA on top of a learning rate schedule, and that the benefits of averaging diminish later in training due to its similarity to learning rate annealing (Sandler et al., 2023). We explore this topic in more details in the following section.
@@
In line with the previous section, we investigate how averaging affects generalization performance across different hyperparameter configurations. This is a common scenario, that may occur when an optimal baseline is not available, or when it is too expensive to search for one. We variate the baseline learning rate and report the achieved validation score with and without averaging. We observe in Figure 6 that the averaged checkpoints closely track the performance of the baseline algorithms, sometimes providing marginal gains. Unlike Sanyal et al. (2023), we do not notice inherent benefits when training at higher learning rates.
@@ 6 Averaging as a Proxy for LR Decay
To achieve optimal performance on most Deep Learning tasks, it is common practice to anneal the learning rate during training (Defazio et al., 2024a). Despite its ubiquity, this practice introduces extra hyperparameters, complicates resuming from the latest checkpoint (Singh et al., 2025) and significantly increases training costs (Hägele et al., 2024). In contrast, averaging weights along the training trajectory intuitively reduces noise and might act as a proxy for learning rate decay (Sandler et al., 2023; Defazio et al., 2024b). Notice, however, that weight averaging also introduces its own hyperparameters that must be tuned (see Appendix B).
@@
Figure 8: Generalization performance when training with NadamW, without ($\bullet$) and with ($\blacktriangle$) learning rate decay, versus training with NadamW +LAWA, without ($\bullet$) and with ($\blacktriangle$) learning rate decay. Despite significantly improving the performance of the baseline, weight averaging alone cannot generally match the validation performance of using a learning rate annealing schedule, and can rarely reach the target validation score. Instead, we observe that combining the two yields slightly better final validation scores throughout the considered workloads. We notice that OGBG and FastMRI are the only workloads where LAWA can provide a good substitute for LR decay, and attribute this behavior to the relatively easy target, very generous step budget, and high variability of these tasks, as also noted by Kasimbeg et al. (2025). We report the mean performance across 3 random seeds. The red dotted line displays the benchmark target scores.
@@
In line with previous work, we compare checkpoint averaging over a long traning run against training with a shorter learning rate schedule. We systematically observe that the validation performance of averaged checkpoints closely tracks a training run with a faster learning rate decay (Figure 7 and Figure 12). This behavior, although pointed out in earlier studies, consistently emerges across the diverse model architectures and datasets considered for this analysis, and explains the considerable speed-ups observed in Section 4. Since checkpoint averaging acts as a proxy for a cooled-down learning rate, WA enables access to a better model during training, thus achieving the target score faster. We note that more sophisticated annealing strategies are possible (Hägele et al., 2024), and while the averaged model consistently approaches the Pareto frontier of loss versus training time (Portes et al., 2022), it may not always lie on it. Nevertheless, its simplicity and effectiveness make averaging a highly practical tool in large-scale training, providing access to a stronger model with minimal computational overhead.
@@
We thoroughly compare averaging with learning rate scheduling across the entire AlgoPerf suite, reporting our findings in Figure 8. Despite heavily tuning the averaging algorithms, we find that fully replacing the learning rate decay with averaging consistently yields inferior performances than annealing the learning rate, with the only exception of Citeo1TB and OGBG, where WA can achieve the target without lowering the learning rate. This result suggests that averaging schemes which do not use the average to compute optimization updates cannot fully replace a learning rate annealing, and more advanced techniques like the one introduced by Defazio et al. (2024b) are necessary to substitute LR decay. Nevertheless, we argue that these averaging schemes can still be used effectively and cheaply to materialize better-performing models during training, as shown in Section 4, with notable applications, including recent work by DeepSeekAI (2025) and Geiping et al. (2025). Additionally, applying them alongside a learning rate decay schedule can further enhance performance and improve generalization, all at minimal cost (Section 5).
@@ 7 Language Modeling
Although the AlgoPerf benchmark covers diverse workloads, including a machine translation task with an encoder–decoder transformer, contemporary language models predominantly adopt decoder-only architectures trained for causal language modeling (next-token prediction). We are therefore interested in understanding and benchmarking the potential impact of weight averaging on this widely adopted application. A previous work (Sanyal et al., 2023) has highlighted the benefits of early weight averaging on modern language models, showing how LAWA can produce stronger models early during training, reaching solid downstream tasks performance faster than the baseline trajectory. Unlike Sanyal et al. (2023), we show that an Exponential Moving Average of model parameters can successfully match the performance gains of LAWA when appropriately tuned. We further investigate the role of averaging hyperparameters, and comment on the connection between sampling frequency, rolling window length, and discount factor in LAWA and EMA.
@@
Since this analysis deviates from AlgoPerf, neither a well-defined target nor a strong baseline are a priori available. In an attempt to fairly evaluate WA on this task, we follow an approach similar to the recent nanogpt speed-runs (Jordan et al., 2024), and set the same target validation loss of $3.28$ as our validation target score. We employ a similar $124$M-parameters model, and the same GPT-2 tokenizer; training on 5B tokens from FineWebEdu (Penedo et al., 2024), and evaluating on a held-out split of $5$M tokens. We use AdamW as the baseline optimizer, and perform a grid search over its hyperparameters to first build a strong reference, as described in Appendix D. We then apply LAWA and EMA on top of the fastest-to-target AdamW configuration and compare against it by means of step-to-target and final validation loss.
@@ Results.
Whereas AdamW reaches the target validation loss in about $3200$ steps at best, both LAWA and EMA can shorten this time by roughly $30\%$ (Figure 9; Figure 14; Table 4). The magnitude of this speed-up clearly dependends on the predefined target, and although the validation loss of the averaged checkpoints is dominated by AdamW’s, the gap between the two shrinks as training proceeds—or more precisely, as the learning rate decreases. Indeed, we observe only minimal benefits in terms of final validation perplexity when comparing the final model obtained with AdamW with the one from LAWA and EMA (Figure 14; Figure 13), similarly to Section 5. Interestingly, we find that, when properly tuned, EMA can match LAWA’s performance while using only a fraction of the memory. Figure 9 displays how EMA hyperparameters affect training speed‐ups, showing how the optimal value of $\beta$ varies with the checkpoint sampling frequency ($\nu$). In our setting, the choice of $\nu=1$, $\beta=0.9$ from Sanyal et al. (2023) is indeed suboptimal, and a higher value of $\beta$ is needed when collecting checkpoint at every step. Similar conclusions hold when evaluating weight averaging by means of final generalization performance, as reported in Appendix D alongside LAWA results.
@@ 8 Limitations
We limit our analysis primarily to NadamW and AdamW as the baseline optimization algorithms. However, we believe that our conclusions are broadly applicable, and demonstrate how averaging can provide substantial benefits also on top of a higher-order optimizer like Shampoo. Nevertheless, it would be interesting to compare averaging schemes across a broader range of optimization algorithms, and against or on top of schedule-free methods (Defazio et al., 2024b).
@@ 9 Conclusions
Through a comprehensive empirical investigation, we study the effects of weight averaging on a challenging optimization benchmark (Dahl et al., 2023), consisting of several diverse Deep Learning architectures and datasets. We evaluate averaging methods by both the time required to reach a reasonable validation target and the best validation performance achieved, concluding that averaging offers substantial efficiency gains while achieving marginally better validation scores. Accessing the optimal model during training without prematurely decaying the learning rate remains an open fascinating challenge, and our work presents weight averaging as a step forward, moving models closer to the Pareto optimality of loss versus training time. We hope that this work further encourages the adoption of averaging techniques. Although such methods are not directly involved in optimizing updates, their significant improvements may dramatically reduce training time. In the field of efficient training, we believe this is a promising study with potential for practical applications (Shen et al., 2024).
@@ A.1 Datasets and Architectures
  • DLRMsmall model on the Criteo 1TB dataset (Naumov et al., 2019; Lab, 2014): a Deep Learning recommendation model optimized for large-scale industrial recommendation tasks.
  • U-Net (Ronneberger et al., 2015) on the FastMRI dataset (Ronneberger et al., 2015; Zbontar et al., 2018): a popular architecture for medical image segmentation.
  • ViT (Vision Transformer) on ImageNet-1k (Dosovitskiy et al., 2020; Deng et al., 2009): a transformer-based model for image classification.
  • GNN model on OGBG dataset (Battaglia et al., 2018; Hu et al., 2020): a graph neural network for learning over graph-structured data, used to test performance on graph-based machine learning tasks.
  • Transformer-big on WMT dataset (Vaswani et al., 2017; Bojar et al., 2017): a large-scale transformer model applied to machine translation.
  • Conformer on LibriSpeech dataset (Gulati et al., 2020; Panayotov et al., 2015): a hybrid architecture combining CNNs and transformers, used for speech recognition tasks.
  • DeepSpeech (Amodei et al., 2016) on LibriSpeech dataset (Panayotov et al., 2015): a recurrent model for speech-to-text, tested on a popular speech recognition dataset to assess its performance in transcribing audio.
@@ A.2 Computational Resources
We make use of the original AlgoPerf repository [^1] https://github.com/mlcommons/algorithmic-efficiency, with minimal modifications, using the PyTorch (Paszke et al., 2017) implementation of the corresponding algorithms. For the Distributed Shampoo implementation, we resort to the original submission to the AlgoPerf competition [^2] https://github.com/mlcommons/algorithms_results_v0.5/tree/main/AlgoPerf_Team_21/external_tuning/shampoo_submission. We select the optimal hyperparameters for NadamW and Shampoo on each workload as the ones providing faster convergence to the benchmark target in the first iteration of the AlgoPerf competition [^3] https://github.com/mlcommons/algorithms_results_v0.5/tree/main/logs/algoperf_scoring_v05/external_tuning/AlgoPerf/prize_qualification_baseline.
@@ Appendix B LAWA and EMA Hyperparameters
Having demonstrated the benefits of averaging weights along the training trajectory, we now investigate the impact of the hyperparameters of the averaging scheme. Specifically, we explore the interaction between the update frequency $\nu$ and the averaging window $L$ for LAWA (Figure 10), as well as the relationship between $\nu$ and $\beta$ for EMA (Figure 11), and investigate how different choices of these hyperparameters influence the effectiveness of averaging in reducing the number of steps required to reach the benchmark validation targets. For this analysis we use NadamW as the underlying optimization algorithm, and the best performing hyperparameter from Dahl et al. (2023) for each workload.
In both cases, we observe a consistent trend: more frequent updates (lower $\nu$) benefit from longer memory (higher $L$ or $\beta$), and vice versa: sampling checkpoints further apart (higher $\nu$) works better with shorter memory buffers (lower $L$ or $\beta$). Several combinations of $\nu$ and $L$ or $\beta$ yield strong performance, suggesting that an optimal averaging horizon might exist. We also note that slower averages (high $L$ or $\beta$) are more sensitive to the choice of $\nu$, requiring more careful tuning, and that EMA is usually more sensitive to its hyperparameters. In practice, we find that values of $\nu$ in the range $64-128$ combined with $L=20$ provide good performance across workloads for LAWA. For EMA, our analysis suggests that a value of $\nu$ between $64$ and $128$ with $\beta=0.9$ is effective. We acknowledge that better configurations may exist, and that the optimal values may vary depending on the workload and on other hyperparameters, particularly on the LR schedule.
@@
Finally, unlike Arpit et al. (2022), we observe that EMA with $\nu=1$ performs poorly for $\beta\leq 0.99$. We find this behavior peculiar, and hypothetize that the worst performance of EMA compared to LAWA observed by Sanyal et al. (2023) might be due to updating the EMA buffer at every step, whereas for LAWA they use larger values of $\nu$, and further explore this in Section 7 and Appendix D. A similar result holds for LAWA, with the only exception of OGBG, where we however acknowledge a very long training horizon and a very slow baseline, as noted in Kasimbeg et al. (2025).
@@ Appendix C Additional AlgoPerf Experiments
We report additional results for the analysis in Section 6, comparing averaging of a long annealed NadamW trajectory with shorter LR schedules across additional workloads. Figure 12 displays the decay schedule and validation performance during training for FastMRI, OGBG, and WMT. On FastMRI and OGBG, both LAWA and EMA closely match the performance of runs with a smaller budget. On OGBG, averaging generally provide better MAP scores throughout training but eventually regress to the overfitting baseline algorithm. On WMT, we observe that a short 25% schedule outperforms the averaging schemes at that point. Despite consistently approaching Pareto optimality of loss versus training time, LAWA and EMA cannot always achieve it. We argue that more sophisticated annealing strategies, possibly similar to Defazio et al. (2024b), may be required to access the optimal model at any time during training.

### v3

@@ 5 Improving Generalization
We report in Table 2 and Figure 8 the results of averaging in terms of generalization performance. We find that using LAWA or EMA on top of a learning rate schedule consistently improves over the baseline optimizer, achieving slightly better validation scores across all the considered workloads. We notably observe significant improvements on the WMT workload. As previously reported in kaddour_stop_2022 and sanyal_early_2023, in the early stage of training, averaging schemes provide considerably better performance than the underlying optimization algorithm, but this gap shrinks towards the end of training. We argue that this behavior is closely linked to the use of WA on top of a learning rate schedule, and that the benefits of averaging diminish later in training due to its similarity to learning rate annealing (sandler_training_2023). We explore this topic in more details in the following section.
@@
In line with the previous section, we investigate how averaging affects generalization performance across different hyperparameter configurations. This is a common scenario, that may occur when an optimal baseline is not available, or when it is too expensive to search for one. We variate the baseline learning rate and report the achieved validation score with and without averaging. We observe in Figure 6 that the averaged checkpoints closely track the performance of the baseline algorithms, sometimes providing marginal gains. Unlike sanyal_early_2023, we do not notice inherent benefits when training at higher learning rates.
@@ 6 Averaging as a Proxy for LR Decay
To achieve optimal performance on most Deep Learning tasks, it is common practice to anneal the learning rate during training (defazio2024optimallineardecaylearning). Despite its ubiquity, this practice introduces extra hyperparameters, complicates resuming from the latest checkpoint (singh2025infinte_schedule) and significantly increases training costs (hagele2024scaling). In contrast, averaging weights along the training trajectory intuitively reduces noise and might act as a proxy for learning rate decay (sandler_training_2023; defazio_road_2024). Notice, however, that weight averaging also introduces its own hyperparameters that must be tuned (see Appendix B).
@@
Figure 8: Generalization performance when training with NadamW, without ($\bullet$) and with ($\blacktriangle$) learning rate decay, versus training with NadamW +LAWA, without ($\bullet$) and with ($\blacktriangle$) learning rate decay. Despite significantly improving the performance of the baseline, weight averaging alone cannot generally match the validation performance of using a learning rate annealing schedule, and can rarely reach the target validation score. Instead, we observe that combining the two yields slightly better final validation scores throughout the considered workloads. We notice that OGBG and FastMRI are the only workloads where LAWA can provide a good substitute for LR decay, and attribute this behavior to the relatively easy target, very generous step budget, and high variability of these tasks, as also noted by kasimbeg2025accelerating. We report the mean performance across 3 random seeds. The red dotted line displays the benchmark target scores.
@@
In line with previous work, we compare checkpoint averaging over a long traning run against training with a shorter learning rate schedule. We systematically observe that the validation performance of averaged checkpoints closely tracks a training run with a faster learning rate decay (Figure 7 and Figure 12). This behavior, although pointed out in earlier studies, consistently emerges across the diverse model architectures and datasets considered for this analysis, and explains the considerable speed-ups observed in Section 4. Since checkpoint averaging acts as a proxy for a cooled-down learning rate, WA enables access to a better model during training, thus achieving the target score faster. We note that more sophisticated annealing strategies are possible (hagele2024scaling), and while the averaged model consistently approaches the Pareto frontier of loss versus training time (portes2022fastbenchmarkingaccuracyvs), it may not always lie on it. Nevertheless, its simplicity and effectiveness make averaging a highly practical tool in large-scale training, providing access to a stronger model with minimal computational overhead.
@@
We thoroughly compare averaging with learning rate scheduling across the entire AlgoPerf suite, reporting our findings in Figure 8. Despite heavily tuning the averaging algorithms, we find that fully replacing the learning rate decay with averaging consistently yields inferior performances than annealing the learning rate, with the only exception of Citeo1TB and OGBG, where WA can achieve the target without lowering the learning rate. This result suggests that averaging schemes which do not use the average to compute optimization updates cannot fully replace a learning rate annealing, and more advanced techniques like the one introduced by defazio_road_2024 are necessary to substitute LR decay. Nevertheless, we argue that these averaging schemes can still be used effectively and cheaply to materialize better-performing models during training, as shown in Section 4, with notable applications, including recent work by DeepSeekV3 and geiping2025scalingtesttimecomputelatent. Additionally, applying them alongside a learning rate decay schedule can further enhance performance and improve generalization, all at minimal cost (Section 5).
@@ 7 Language Modeling
Although the AlgoPerf benchmark covers diverse workloads, including a machine translation task with an encoder–decoder transformer, contemporary language models predominantly adopt decoder-only architectures trained for causal language modeling (next-token prediction). We are therefore interested in understanding and benchmarking the potential impact of weight averaging on this widely adopted application. A previous work (sanyal_early_2023) has highlighted the benefits of early weight averaging on modern language models, showing how LAWA can produce stronger models early during training, reaching solid downstream tasks performance faster than the baseline trajectory. Unlike sanyal_early_2023, we show that an Exponential Moving Average of model parameters can successfully match the performance gains of LAWA when appropriately tuned. We further investigate the role of averaging hyperparameters, and comment on the connection between sampling frequency, rolling window length, and discount factor in LAWA and EMA.
@@
Since this analysis deviates from AlgoPerf, neither a well-defined target nor a strong baseline are a priori available. In an attempt to fairly evaluate WA on this task, we follow an approach similar to the recent nanogpt speed-runs (modded_nanogpt_2024), and set the same target validation loss of $3.28$ as our validation target score. We employ a similar 124124M-parameters model, and the same GPT-2 tokenizer; training on 5B tokens from FineWebEdu (penedo2024finewebedu), and evaluating on a held-out split of 55M tokens. We use AdamW as the baseline optimizer, and perform a grid search over its hyperparameters to first build a strong reference, as described in Appendix D. We then apply LAWA and EMA on top of the fastest-to-target AdamW configuration and compare against it by means of step-to-target and final validation loss.
@@ Results.
Whereas AdamW reaches the target validation loss in about $3200$ steps at best, both LAWA and EMA can shorten this time by roughly $30\%$ (Figure 9; Figure 14; Table 4). The magnitude of this speed-up clearly dependends on the predefined target, and although the validation loss of the averaged checkpoints is dominated by AdamW’s, the gap between the two shrinks as training proceeds—or more precisely, as the learning rate decreases. Indeed, we observe only minimal benefits in terms of final validation perplexity when comparing the final model obtained with AdamW with the one from LAWA and EMA (Figure 14; Figure 13), similarly to Section 5. Interestingly, we find that, when properly tuned, EMA can match LAWA’s performance while using only a fraction of the memory. Figure 9 displays how EMA hyperparameters affect training speed‐ups, showing how the optimal value of $\beta$ varies with the checkpoint sampling frequency ($\nu$). In our setting, the choice of $\nu=1$, $\beta=0.9$ from sanyal_early_2023 is indeed suboptimal, and a higher value of $\beta$ is needed when collecting checkpoint at every step. Similar conclusions hold when evaluating weight averaging by means of final generalization performance, as reported in Appendix D alongside LAWA results.
@@ 8 Limitations
We limit our analysis primarily to NadamW and AdamW as the baseline optimization algorithms. However, we believe that our conclusions are broadly applicable, and demonstrate how averaging can provide substantial benefits also on top of a higher-order optimizer like Shampoo. Nevertheless, it would be interesting to compare averaging schemes across a broader range of optimization algorithms, and against or on top of schedule-free methods (defazio_road_2024).
@@ 9 Conclusions
Through a comprehensive empirical investigation, we study the effects of weight averaging on a challenging optimization benchmark (dahl_benchmarking_2023), consisting of several diverse Deep Learning architectures and datasets. We evaluate averaging methods by both the time required to reach a reasonable validation target and the best validation performance achieved, concluding that averaging offers substantial efficiency gains while achieving marginally better validation scores. Accessing the optimal model during training without prematurely decaying the learning rate remains an open fascinating challenge, and our work presents weight averaging as a step forward, moving models closer to the Pareto optimality of loss versus training time. We hope that this work further encourages the adoption of averaging techniques. Although such methods are not directly involved in optimizing updates, their significant improvements may dramatically reduce training time. In the field of efficient training, we believe this is a promising study with potential for practical applications (shen_efficency).
@@ A.1 Datasets and Architectures
  • DLRMsmall model on the Criteo 1TB dataset (naumov2019deep; criteo2014dataset): a Deep Learning recommendation model optimized for large-scale industrial recommendation tasks.
  • U-Net (unet_2015) on the FastMRI dataset (unet_2015; fastMRI_dataset): a popular architecture for medical image segmentation.
  • ViT (Vision Transformer) on ImageNet-1k (dosovitskiy2020image; deng2009imagenet): a transformer-based model for image classification.
  • GNN model on OGBG dataset (battaglia2018relational; hu2020open): a graph neural network for learning over graph-structured data, used to test performance on graph-based machine learning tasks.
  • Transformer-big on WMT dataset (vaswani2017attention; bojar2017findings): a large-scale transformer model applied to machine translation.
  • Conformer on LibriSpeech dataset (gulati2020conformer; librispeech): a hybrid architecture combining CNNs and transformers, used for speech recognition tasks.
  • DeepSpeech (deepspeech_amodei16) on LibriSpeech dataset (librispeech): a recurrent model for speech-to-text, tested on a popular speech recognition dataset to assess its performance in transcribing audio.
@@ A.2 Computational Resources
We make use of the original AlgoPerf repository [^1] https://github.com/mlcommons/algorithmic-efficiency, with minimal modifications, using the PyTorch (paszke2017pytorch) implementation of the corresponding algorithms. For the Distributed Shampoo implementation, we resort to the original submission to the AlgoPerf competition [^2] https://github.com/mlcommons/algorithms_results_v0.5/tree/main/AlgoPerf_Team_21/external_tuning/shampoo_submission. We select the optimal hyperparameters for NadamW and Shampoo on each workload as the ones providing faster convergence to the benchmark target in the first iteration of the AlgoPerf competition [^3] https://github.com/mlcommons/algorithms_results_v0.5/tree/main/logs/algoperf_scoring_v05/external_tuning/AlgoPerf/prize_qualification_baseline.
@@ Appendix B LAWA and EMA Hyperparameters
Having demonstrated the benefits of averaging weights along the training trajectory, we now investigate the impact of the hyperparameters of the averaging scheme. Specifically, we explore the interaction between the update frequency $\nu$ and the averaging window $L$ for LAWA (Figure 10), as well as the relationship between $\nu$ and $\beta$ for EMA (Figure 11), and investigate how different choices of these hyperparameters influence the effectiveness of averaging in reducing the number of steps required to reach the benchmark validation targets. For this analysis we use NadamW as the underlying optimization algorithm, and the best performing hyperparameter from dahl_benchmarking_2023 for each workload.
In both cases, we observe a consistent trend: more frequent updates (lower $\nu$) benefit from longer memory (higher $L$ or $\beta$), and vice versa: sampling checkpoints further apart (higher $\nu$) works better with shorter memory buffers (lower $L$ or $\beta$). Several combinations of $\nu$ and $L$ or $\beta$ yield strong performance, suggesting that an optimal averaging horizon might exist. We also note that slower averages (high $L$ or $\beta$) are more sensitive to the choice of $\nu$, requiring more careful tuning, and that EMA is usually more sensitive to its hyperparameters. In practice, we find that values of $\nu$ in the range 6412864-128 combined with $L=20$ provide good performance across workloads for LAWA. For EMA, our analysis suggests that a value of $\nu$ between $64$ and $128$ with $\beta=0.9$ is effective. We acknowledge that better configurations may exist, and that the optimal values may vary depending on the workload and on other hyperparameters, particularly on the LR schedule.
@@
Finally, unlike arpit2022ensembleaveragesimprovingmodel, we observe that EMA with $\nu=1$ performs poorly for $\beta\leq 0.99$. We find this behavior peculiar, and hypothetize that the worst performance of EMA compared to LAWA observed by sanyal_early_2023 might be due to updating the EMA buffer at every step, whereas for LAWA they use larger values of $\nu$, and further explore this in Section 7 and Appendix D. A similar result holds for LAWA, with the only exception of OGBG, where we however acknowledge a very long training horizon and a very slow baseline, as noted in kasimbeg2025accelerating.
@@ Appendix C Additional AlgoPerf Experiments
We report additional results for the analysis in Section 6, comparing averaging of a long annealed NadamW trajectory with shorter LR schedules across additional workloads. Figure 12 displays the decay schedule and validation performance during training for FastMRI, OGBG, and WMT. On FastMRI and OGBG, both LAWA and EMA closely match the performance of runs with a smaller budget. On OGBG, averaging generally provide better MAP scores throughout training but eventually regress to the overfitting baseline algorithm. On WMT, we observe that a short 25% schedule outperforms the averaging schemes at that point. Despite consistently approaching Pareto optimality of loss versus training time, LAWA and EMA cannot always achieve it. We argue that more sophisticated annealing strategies, possibly similar to defazio_road_2024, may be required to access the optimal model at any time during training.